OpenStudy (anonymous):

Can someone help me verify a trig identity. (cos x / 1+sin x)+(1+sin x/cos x)=2 sec x

OpenStudy (anonymous):

I've got it to (cos x / 1+sin x)+(1+sin x/cos x)=1/cos x, but I don't think I'm doing it right.

OpenStudy (anonymous):

@AllTehMaffs

OpenStudy (trojanpoem):

cosx/1+sinx + 1+sinx/cosx = cos^2x + (1+sinx)^2 / cosx(1+ sinx) cos^2x + sin^2x + 1 + 2sinx / cosx(1+sinx) (Note :cos^2x + sin^2x = 1) 2(1+sinx)/cosx(1+sinx) = 2/cosx = 2 secx

OpenStudy (anonymous):

Oh ok I thought I was doing it wrong thanks alot dude.

OpenStudy (trojanpoem):

yw